Presented below is information related to Monty Manufacturing Corporation Estimated Salvage Estimated Life (in years) Asset Cost $47,790 39,648 42,480 22,420 27,730 $6,490 5,664 4,248 1,770 2,950 10 Compute the rate of depreciation per year to be applied to the plant assets under the composite method. (Round answer to 2 decimal place, eg, 4.83%.) Composite rate SHOW LIST OF ACCOUNTS LINK TO TEXT LINK TO TEXT Prepare the adjusting entry necessary at the end of the year to record depreciation for the year. (Ifno entry is required, select No entry for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually. Use Plant Assets related account.) Account Titles and Explanation Debit CreditPrepare the entry to record the sale of asset D for cash of牛5,664. It was used for 6 years, and depreciation was entered under the composite method. (If no entry is required, select No entry for the account titles and enter 0 for the amour Credit account titles are automatically indented when amount is entered. Do not indent manually. Use Plant Assets related account.) Account Titles and Explanation Credit
Answer #1

SOLUTION

1. Depreciation expense = (Cost of assets - Salvage value) / Estimated useful life of asset

Asset A-

Depreciation expense = ($47,790 - $6,490) / 10 years = $4,130

Asset B-

Depreciation expense = ($39,648 - $5,664) / 9 years = $3,776

Asset C-

Depreciation expense = ($42,480 - $4,248) / 9 years = $4,248

Asset D-

Depreciation expense = ($22,420 - $1,770) / 7 years = $2,950

Asset E-

Depreciation expense = ($27,730 - $2,950) / 6 years = $4,130

Total cost of the assets = Cost of asset A + Cost of asset B + Cost of asset C + Cost of asset D + Cost of asset E

= $47,790+$39,648+$42,480+$22,420+$27,730 = 180,068

Total depreciation = $4,130+$3,776+$4,248+$2,950+$4,130 = 19,234

Rate of depreciation = Total depreciation / Total cost of the assets

= $19,234 / $180,068 = 10.68%

2. Journal entry-

S.No. Accounts titles and Explanation Debit ($) Credit ($)
1. Depreciation expense 19,234
Accumulated depreciation 19,234
(To record the depreciation expense)

Depreciation expense = Total cost of the assets * Rate of depreciation

= $180,068 * 10.68% = $19,234

3. Journal entry-

S.No. Accounts titles and Explanation Debit ($) Credit ($)
1. Cash 5,664
Accumulated depreciation 14,367
Loss on sales 2,389
Asset D 22,420
(To record the depreciation expense)

Accumulated depreciation = (22,420*10.68%) * 6 years

= 14,367